Suppose a coin is tossed 6 times, find the probability of obtaining between 4 and 6 heads inclusively.
Answer by stanbon(75887)


Probability-and-statistics/1079639 (2017-05-04 12:58:49): Suppose a coin is tossed 6 times, find the probability of obtaining between 4 and 6 heads inclusively.
--------------------
Binomial Problem with n = 6 and p(head) = 1/2
--------
P(4<= x <= 6) = 1 - P(0<= x <=3) = 1- binomcdf(6,.0.5,3) = 0.3438
------------
Cheers,
Stan H.
-------------